OpenStudy (anonymous):

Can somebody please help me CREATE these functions, I am retarded when it comes to making my own equations. One function, f(x), with two real rational solutions. One function, g(x), with two real irrational solutions. One function, h(x), with two complex solutions.

OpenStudy (anonymous):

This is what I have going off some of the functions in my lesson, but I'm not sure if there right. \[f(x)=(x-3)(x-2)\] \[g(x)=(x-\sqrt{2})(x-\sqrt{3})\] \[h(x)=(x-2i)(x-3i)\]

OpenStudy (anonymous):

@RC_CA @mathstudent55